OpenStudy (anonymous):

Which of the following is a solution of x2 - 2x = -8

OpenStudy (anonymous):

can u explain

geerky42 (geerky42):

Get it in form of \(ax^2+bx+c =0\), then either try to factor it, or just use quadratic formula.

OpenStudy (anonymous):

so its x2-2x-8=0

geerky42 (geerky42):

well, you get it to form it need to be. now you have to find the value of x.

geerky42 (geerky42):

Try to factor it

OpenStudy (solomonzelman):

\[x^2-2x=8\]\[x^2-2x+1=8+1\]\[(x-1)^2=9\]\[x-1=±~3\]\[x=1~±~3~~~~~~~~~~~~--->~~~~~~~~~~~x=~~4,~~~-2\]

OpenStudy (solomonzelman):

The beauty of completing the square (well I skipped little steps )

OpenStudy (anonymous):

negative 1 minus i square root of 28 1 plus i square root of 28 negative 1 minus i square root of 7 1 plus i square root of 7

OpenStudy (anonymous):

i dont understand

geerky42 (geerky42):

I think these options are for other question. SolomonZelman is correct.